# Accelerating Diffusion Transformers with Token-wise Feature Caching

### v2

@@ 2 Related Work
Transformers in Diffusion Models Diffusion models (DMs) (Ho et al., 2020; Sohl-Dickstein et al., 2015), which iteratively denoise an initial noise input through a series of diffusion steps, have achieved remarkable success across various generation applications (Rombach et al., 2022; Balaji et al., 2022). Early DMs (Ho et al., 2020; Rombach et al., 2022) are based on the U-Net architecture (Ronneberger et al., 2015), consistently achieving satisfactory generation results. Recently, Diffusion Transformer (DiT) (Peebles & Xie, 2023) has emerged as a major advancement by replacing the U-Net backbone with a Transformer architecture. This transition enhances the scalability and efficiency of DMs across various generative tasks (Chen et al., 2024a; Brooks et al., 2024). For example, PixArt-$\alpha$ (Chen et al., 2024a) utilizes DiT as a scalable foundational model, adapting it for text-to-image generation, while Sora (Brooks et al., 2024) demonstrates DiT’s potential in high-fidelity video generation, inspiring a series of related open-source projects (Zheng et al., 2024; Lab & etc., 2024). Despite their success, the iterative denoising process of these DMs is significantly time-consuming, making them less feasible for practical applications.
@@
Acceleration of Diffusion Models To improve the generation efficiency of DMs, numerous diffusion acceleration methods have been proposed, falling broadly into two categories: (1) reducing the number of sampling timesteps, and (2) accelerating the denoising networks. The first category aims to achieve high-quality generation results with fewer sampling steps. DDIM (Song et al., 2021) introduces a deterministic sampling process that reduces the number of denoising steps while preserving generation quality. DPM-Solver (Lu et al., 2022a) and DPM-Solver++ (Lu et al., 2022b) propose adaptive high-order solvers for a faster generation without compromising on generation results. Rectified flow (Liu et al., 2023) optimizes distribution transport in ODE models to facilitate efficient and high-quality generation, enabling sampling with fewer timesteps. In the second category, various efforts have been paid to token reduction (Bolya & Hoffman, 2023), knowledge distillation (Li et al., 2024), and weight quantization (Shang et al., 2023) and pruning (Fang et al., 2023) on the denoising networks. Additionally, recent cache-based methods reduce redundant computations to accelerate inference for DMs. These cache-based methods have obtained abundant attention since they have no requirements for additional training. DeepCache (Ma et al., 2024b) eliminates redundant computations in Stable Diffusion (Rombach et al., 2022) by reusing intermediate features of low-resolution layers in the U-Net. Faster Diffusion (Li et al., 2023) accelerates the sampling process of DMs by caching U-Net encoder features across timesteps, skipping encoder computations at certain steps. Unfortunately, DeepCache and Faster Diffusion are designed specifically for U-Net-based denoisers and can not be applied to DiT (Chen et al., 2024b). Recently, FORA (Selvaraju et al., 2024) and $\Delta$-DiT (Chen et al., 2024b) have been proposed to cache the features and the residual of features for DiT. Learning-to-Cache (Ma et al., 2024a) learns an optimal cache strategy, which achieves a slightly higher acceleration ratio but introduces the requirements of training. However, these methods apply the identical cache solution to all the tokens and even all the layers, which leads to a significant performance degradation in generation quality.
@@ 3.5 Token Selection
(IV) Uniform Spatial Distribution: The pixels in the neighboring patch of the images usually contain similar information. As discussed in previous works, overwhelmingly pruning the information in a local spatial region may result in significant performance degradation in the whole images (Bolya & Hoffman, 2023). Hence, to guarantee that the errors introduced by caching are not densely distributed in the same spatial region, we define the following scoring function: $s_{4}(x_{i})=\mathcal{I}(x_{i})\cdot\left(s_{1}(x_{i})+s_{2}(x_{i})+s_{3}(x_{i})\right),$ where $\mathcal{I}(x_{i})$ is an indicator function which equals to 1 if $x_{i}$ has the highest score of $s_{1}(x_{i})+s_{2}(x_{i})+s_{3}(x_{i})$ in its neighboring $k\times k$ pixels and equals to 0 in the other settings.
@@
$\mathcal{I}_{~{}\text{Cache}}=\mathop{\arg\min}_{\{i_{1},i_{2},\dots,i_{R\%\times N}\}\subseteq{\{1,2,\dots,n}\}}\{\mathcal{S}(x_{i_{1}}),\mathcal{S}(x_{i_{2}}),\cdots,\mathcal{S}(x_{i_{n}})\}.$ (3)
@@ 4.1 Experiment Settings
Model Configurations We conduct experiments on three commonly-used DiT-based models across different generation tasks, including PixArt-$\alpha$ (Chen et al., 2024a) for text-to-image generation, OpenSora (Zheng et al., 2024) for text-to-video generation, and DiT-XL/2 (Peebles & Xie, 2023) for class-conditional image generation with NVIDIA A800 80GB GPUs. Each model utilizes its default sampling method: DPM-Solver++ (Lu et al., 2022b) with 20 steps for PixArt-$\alpha$, rflow (Liu et al., 2023) with 30 steps for OpenSora and DDIM (Song et al., 2021) with 250 steps for DiT-XL/2. For each model, we configure different average forced activation cycles $\mathcal{N}$ and average caching ratios $R$ for ToCa as follows: PixArt-$\alpha$: $\mathcal{N}=3$ and $R=70\%$, OpenSora: $\mathcal{N}=3$ for temporal attention, spatial attention, MLP, and $\mathcal{N}=6$ for cross-attention, with $R=85\%$ exclusively for MLP, and DiT: $\mathcal{N}=4$ and $R=93\%$. Please refer to the appendix for more implementation details.
@@ 4.2 Results on Text-to-Image Generation
In Table 1, we compare ToCa configured with parameters to achieve an acceleration ratio close to 2.0, against two other training-free acceleration approaches: FORA (Selvaraju et al., 2024), a recent cache-based high-acceleration method, and the 10-step DPM-Solver++ sampling (Lu et al., 2022b). In terms of generation quality, the quantitative results demonstrate that ToCa achieves the lowest FID among the compared acceleration methods while maintaining a high acceleration ratio. Figure LABEL:fig:PixArt-vis also illustrates that our generated results most closely resemble those of the original PixArt-$\alpha$. Regarding generation consistency, Table 1 demonstrates that ToCa achieves the highest CLIP score on both MS-COCO2017 and the PartiPrompts. Figure LABEL:fig:PixArt-vis shows that ToCa generates images that align more closely with the text descriptions compared to other methods. This is particularly evident in the fourth case, where only ToCa successfully generates an image matching ”a small rowboat tied to a wooden dock”, while other methods fail to generate the content of ”a wooden dock”. This may be caused by cross-attention score $s_{2}$ in ToCa that ensures the frequent refreshing of tokens that are semantically relevant to the text descriptions, resulting in generated images with enhanced semantic consistency to the text prompts.
@@ 4.3 Results on Text-to-Video Generation
We compare ToCa with adjusted rflow sampling steps from 20 to 10, alongside other acceleration methods including FORA, PAB (Zhao et al., 2024), $\Delta$-DiT (Chen et al., 2024b), and T-GATE (Zhang et al., 2024) using OpenSora (Zheng et al., 2024) for text-to-video generation. As presented in Table 2, the experimental results show that ToCa achieves an impressive VBench score offering the lowest computational cost and highest inference speed among all methods tested. The VBench score of the 2.36×2.36$\times$2.36 × accelerated ToCa scheme drops by only 0.79 compared to the non-accelerated scheme, while FORA’s score decreases by 2.22, resulting in a $64.4\%$ reduction in quality loss. Additionally, more VBench metrics results are presented in Figure 8, which illustrate that ToCa significantly speeds up the original OpenSora with only slight performance degradation on a few metrics. Notably, ToCa stands out as the sole acceleration method achieving nearly overall consistency performance with the original OpenSora, clearly outperforming another cache-based acceleration method FORA. This again highlights the effectiveness of our proposed cross-attention-based token selection strategy, ensuring that the generated videos are highly aligned with the text descriptions. We further present some video generation results in Figure 7, where we observe that the visual fidelity, and overall consistency of ToCa are closest to the original OpenSora. Please kindly refer to our video demos in the supplementary material or the anonymous web page for viewing the video demo.
@@ 4.4 Results on Class-Conditional Image Generation
DiT-XL/2-G (cfg $=1.50$) 2.012 59.36 1.00$\times$ 4.98 2.31 0.82 0.58
$33\%$ steps 0.681 19.71 3.01$\times$ 6.31 2.76 0.81 0.57
$37\%$ steps 0.748 22.08 2.69$\times$ 6.04 2.64 0.81 0.58
$\textbf{FORA}(\mathcal{N}=3)$ 0.807 20.02 2.97$\times$ 6.21 2.80 0.80 0.59
$\textbf{FORA}(\mathcal{N}=2.8)$ 0.815 21.68 2.74$\times$ 6.13 2.80 0.80 0.59
ToCa ($\mathcal{N}=4,R=93\%$) 0.820 21.61 2.75$\times$ 5.74 2.58 0.81 0.59

### v4

@@ 2 Related Work
Diffusion models (DMs) (Ho et al., 2020; Sohl-Dickstein et al., 2015), which iteratively denoise an initial noise input through a series of diffusion steps, have achieved remarkable success across various generation applications (Rombach et al., 2022; Balaji et al., 2022). Early DMs (Ho et al., 2020; Rombach et al., 2022) are based on the U-Net architecture (Ronneberger et al., 2015), consistently achieving satisfactory generation results. Recently, Diffusion Transformer (DiT) (Peebles & Xie, 2023) has emerged as a major advancement by replacing the U-Net backbone with a Transformer architecture. This transition enhances the scalability and efficiency of DMs across various generative tasks (Chen et al., 2024a; Brooks et al., 2024). For example, PixArt-$\alpha$ (Chen et al., 2024a) utilizes DiT as a scalable foundational model, adapting it for text-to-image generation, while Sora (Brooks et al., 2024) demonstrates DiT’s potential in high-fidelity video generation, inspiring a series of related open-source projects (Zheng et al., 2024; Lab & etc., 2024). Despite their success, the iterative denoising process of these DMs is significantly time-consuming, making them less feasible for practical applications.
Acceleration of Diffusion Models
@@
To improve the generation efficiency of DMs, numerous diffusion acceleration methods have been proposed, falling broadly into two categories: (1) reducing the number of sampling timesteps, and (2) accelerating the denoising networks. The first category aims to achieve high-quality generation results with fewer sampling steps. DDIM (Song et al., 2021) introduces a deterministic sampling process that reduces the number of denoising steps while preserving generation quality. DPM-Solver (Lu et al., 2022a) and DPM-Solver++ (Lu et al., 2022b) propose adaptive high-order solvers for a faster generation without compromising on generation results. Rectified flow (Liu et al., 2023) optimizes distribution transport in ODE models to facilitate efficient and high-quality generation, enabling sampling with fewer timesteps. Step-distillation (Salimans & Ho, 2022; Meng et al., 2023) minimizes the number of timesteps with knowledge distillation from multiple timesteps to fewer ones. Consistency models (Song et al., 2023) accelerate generative modeling by mapping noise directly to data and enforcing self-consistency across steps. In the second category, various efforts have been paid to token reduction (Bolya & Hoffman, 2023; Zhang et al., 2025; 2024a), knowledge distillation (Li et al., 2024), and weight quantization (Li et al., 2023b; Shang et al., 2023) and pruning (Fang et al., 2023) on the denoising networks. Additionally, recent cache-based methods reduce redundant computations to accelerate inference for DMs. These cache-based methods have obtained abundant attention since they have no requirements for additional training. DeepCache (Ma et al., 2024b) eliminates redundant computations in Stable Diffusion (Rombach et al., 2022) by reusing intermediate features of low-resolution layers in the U-Net. Faster Diffusion (Li et al., 2023a) accelerates the sampling process of DMs by caching U-Net encoder features across timesteps, skipping encoder computations at certain steps. Unfortunately, DeepCache and Faster Diffusion are designed specifically for U-Net-based denoisers and can not be applied to DiT (Chen et al., 2024b). Recently, FORA (Selvaraju et al., 2024) and $\Delta$-DiT (Chen et al., 2024b) have been proposed to cache the features and the residual of features for DiT. Learning-to-Cache (Ma et al., 2024a) learns an optimal cache strategy, which achieves a slightly higher acceleration ratio but introduces the requirements of training. However, these methods apply the identical cache solution to all the tokens and even all the layers, which leads to a significant performance degradation in generation quality.
@@ 3 Methodology
In the Methodology section, we briefly introduce Diffusion Models and the feature caching acceleration method for Diffusion Transformers, followed by the ToCa workflow and important token selection. The importance of a token $x_{i}$ is determined by $s_{1}$ interaction strength with other tokens, $s_{2}$ association with global textual information, $s_{3}$ accumulated cache error, which, if excessive, can cause image collapse and should be controlled, and $s_{4}$ spatial uniformity of selected tokens.
@@ 3.5 Token Selection
(IV) Uniform Spatial Distribution: The pixels in the neighboring patch of the images usually contain similar information. As discussed in previous works, overwhelmingly pruning the information in a local spatial region may result in significant performance degradation in the whole images (Bolya & Hoffman, 2023). Hence, to guarantee that the errors introduced by caching are not densely distributed in the same spatial region, we define the following scoring function: $s_{4}(x_{i})=\mathcal{I}(x_{i})\cdot\left(\lambda_{1}\cdot s_{1}(x_{i})+\lambda_{2}\cdot s_{2}(x_{i})+\lambda_{3}\cdot s_{3}(x_{i})\right),$ where $\mathcal{I}(x_{i})$ is an indicator function which equals to 1 if $x_{i}$ has the highest score of $\lambda_{1}\cdot s_{1}(x_{i})+\lambda_{2}\cdot s_{2}(x_{i})+\lambda_{3}\cdot s_{3}(x_{i})$ in its neighboring $k\times k$ pixels and 0 in the other settings, and $\lambda_{j}$ are hyper-parameters to balance each score.
@@
$\mathcal{I}_{~{}\text{Cache}}=\mathop{\arg\min}_{\{i_{1},i_{2},\dots,i_{R\%\times N}\}\subseteq{\{1,2,\dots,n}\}}\{\mathcal{S}(x_{i_{1}}),\mathcal{S}(x_{i_{2}}),\cdots,\mathcal{S}(x_{i_{n}})\}.\vspace{-0.4cm}$ (3)
@@ 4.1 Experiment Settings
Model Configurations We conduct experiments on three commonly-used DiT-based models across different generation tasks, including PixArt-$\alpha$ (Chen et al., 2024a) for text-to-image generation, OpenSora (Zheng et al., 2024) for text-to-video generation, and DiT-XL/2 (Peebles & Xie, 2023) for class-conditional image generation with NVIDIA A800 80GB GPUs. Each model utilizes its default sampling method: DPM-Solver++ (Lu et al., 2022b) with 20 steps for PixArt-$\alpha$, rflow (Liu et al., 2023) with 30 steps for OpenSora and DDPM (Ho et al., 2020) with 250 steps for DiT-XL/2. For each model, we configure different average forced activation cycles $\mathcal{N}$ and average caching ratios $R$ for ToCa as follows: PixArt-$\alpha$: $\mathcal{N}=3$ and $R=70\%$, OpenSora: $\mathcal{N}=3$ for temporal attention, spatial attention, MLP, and $\mathcal{N}=6$ for cross-attention, with $R=85\%$ exclusively for MLP, and DiT: $\mathcal{N}=4$ and $R=93\%$. Please refer to the appendix for more implementation details.
@@ 4.2 Results on Text-to-Image Generation
In Table 1, we compare ToCa configured with parameters to achieve an acceleration ratio close to 2.0, against two other training-free acceleration approaches: FORA (Selvaraju et al., 2024), a recent cache-based high-acceleration method, and the 10-step DPM-Solver++ sampling (Lu et al., 2022b). In terms of generation quality, the quantitative results demonstrate that ToCa achieves the lowest FID among the compared acceleration methods while maintaining a high acceleration ratio. Figure LABEL:fig:PixArt-vis also illustrates that our generated results most closely resemble those of the original PixArt-$\alpha$. Regarding generation consistency, Table 1 demonstrates that ToCa achieves the highest CLIP score on both MS-COCO2017(Lin et al., 2014) and the PartiPrompts(Yu et al., 2022). Figure LABEL:fig:PixArt-vis shows that ToCa generates images that align more closely with the text descriptions compared to other methods. This is particularly evident in the fourth case, where only ToCa successfully generates an image matching ”a small rowboat tied to a wooden dock”, while other methods fail to generate the content of ”a wooden dock”. This may be caused by cross-attention score $s_{2}$ in ToCa that ensures the frequent refreshing of tokens that are semantically relevant to the text descriptions, resulting in generated images with enhanced semantic consistency to the text prompts.
@@ 4.3 Results on Text-to-Video Generation
We compare ToCa with adjusted rflow sampling steps from 20 to 10, alongside other acceleration methods including FORA, PAB (Zhao et al., 2024), $\Delta$-DiT (Chen et al., 2024b), and T-GATE (Zhang et al., 2024b) using OpenSora (Zheng et al., 2024) for text-to-video generation. As presented in Table 2, the experimental results show that ToCa achieves an impressive VBench score offering the lowest computational cost and highest inference speed among all methods tested. The VBench score of the 2.36×2.36$\times$2.36 × accelerated ToCa scheme drops by only 0.79 compared to the non-accelerated scheme, while FORA’s score decreases by 2.22, resulting in a $64.4\%$ reduction in quality loss. Additionally, more VBench metrics results are presented in Figure 8, which illustrate that ToCa significantly speeds up the original OpenSora with only slight performance degradation on a few metrics. Notably, ToCa stands out as the sole acceleration method achieving nearly overall consistency performance with the original OpenSora, clearly outperforming another cache-based acceleration method FORA. This again highlights the effectiveness of our proposed cross-attention-based token selection strategy, ensuring that the generated videos are highly aligned with the text descriptions. We further present some video generation results in Figure 7, where we observe that the visual fidelity, and overall consistency of ToCa are closest to the original OpenSora. Please kindly refer to our video demos in the supplementary material or the anonymous web page for viewing the video demo.
@@ 4.4 Results on Class-Conditional Image Generation
DiT-XL/2-G (cfg $=1.50$) 2.012 118.68 1.00$\times$ 4.98 2.31 0.82 0.58
$33\%$ steps 0.681 39.40 3.01$\times$ 6.31 2.76 0.81 0.57
$37\%$ steps 0.748 44.15 2.69$\times$ 6.04 2.64 0.81 0.58
$\textbf{FORA}(\mathcal{N}=3)$ 0.807 39.95 2.97$\times$ 6.21 2.80 0.80 0.59
$\textbf{FORA}(\mathcal{N}=2.8)$ 0.815 43.36 2.74$\times$ 6.13 2.80 0.80 0.59
ToCa ($\mathcal{N}=4,R=93\%$) 0.820 43.22 2.75$\times$ 5.74 2.58 0.81 0.59
@@ A.3 Sensitivity Study
A.4 Computation Complexity Analysis
A.4.1 Main Computations
Complexity of Attention Layer.
In the Attention layer, tokens are first processed through a linear layer to generate queries, keys, and values. Next, the queries and keys are multiplied using a dot product, passed through a softmax function, and then multiplied with the values. Finally, the result is passed through another linear layer to produce the output. The computational cost of the Self-Attention layer can be expressed as:
$\text{FLOPs}_{SA}\approx$ $N\times D\times 3\times D\times 2+N^{2}\times D\times 2+N^{2}\times H\times 5+N^{2}\times D\times 2+N\times D\times D\times 2$ (4)
$=$ $8\times N\times D^{2}+4\times N^{2}\times D+5\times N^{2}\times H\approx O(ND^{2})+O(N^{2}D),$
where $N$ is the number of tokens, $D$ is the hidden state of each token and $H$ is the number of heads, ($H\ll D$). The computational cost of a Cross-Attention layer can be expressed as:
$\text{FLOPs}_{CA}\approx$ $N_{1}\times D\times D\times 2+N_{2}\times D\times 2\times D\times 2+N_{1}\times N_{2}\times D\times 2$ (5)
$+N_{1}\times N_{2}\times H\times 5+N_{1}\times N_{2}\times D\times 2+N_{1}\times D\times D\times 2$
$=$ $4\times(N_{1}+N_{2})\times D^{2}+4\times N_{1}\times N_{2}\times D+5\times N_{1}\times N_{2}\times H$
$\approx$ $O((N_{1}+N_{2})D^{2})+O(N_{1}N_{2}D)=O((N+N_{2})D^{2})+O(NN_{2}D),$
where $N_{1}=N,N_{2}$ are the number of image and text tokens, D is the hidden state of each token and H is the number of heads, ($H\ll D$). In the previous computations, the softmax operation is approximated as involving 5 floating-point calculations per element.
Complexity of MLP Layer.
The computational cost of MLP layer can be written as:
$\text{FLOPs}_{MLP}\approx$ $N\times D_{1}\times D_{2}\times 2+N\times D_{2}\times 6+N\times D_{1}\times D_{2}\times 2$ (6)
$=$ $4\times N\times D_{1}\times D_{2}+6\times D_{2}\times N$
$=$ $16\times N\times D^{2}+24\times N\times D\approx O(ND^{2}),$
where $N$ is the number of tokens, $D_{1}=D$ and $D_{2}=4D_{1}$ are the hidden and middle-hidden state in MLP, respectively. The activation function for MLP is approximated to involve 6 floating-point operations per element.
A.4.2 Computation costs from token selection
Self-Attention score $s_{1}$.
As mentioned in section 3.5, the Self-Attention score $s_{1}$ is computed as $s_{1}(x_{i})=\sum_{i=1}^{N}\alpha_{ij}$, where $x_{i}$ is the $i_{th}$ token, and $\alpha_{ij}$ is the element in the self-attention map. Therefore, the computational complexity of the self-attention score is only $N\approx O(N)$. In a practical case achieving about 2.3×2.3$\times$2.3 × acceleration, the computation cost of the Self-Attention score accounts for approximately $0.28\%$ of the main components.
Cross-Attention score $s_{2}$.
As mentioned in section 3.5, the Cross-Attention score $s_{2}$ is computed as $s_{2}(x_{i})=-\sum_{j=1}^{N}c_{ij}\log(c_{ij})$, where the $c_{ij}$ is the element in the cross-attention map. Therefore, the computational complexity of the cross-attention score is only $2N\approx O(N)$. In a practical case achieving about 2.3×2.3$\times$2.3 × acceleration, the computation cost of the Cross-Attention score accounts for approximately $0.35\%$ of the main components.
Cache Frequency score $s_{3}$ and Uniform Spatial Distribution score $s_{4}$.
[…]
In summary, although ToCa introduces additional computations, its computational complexity of $O(N)$ is negligible compared to the main computational modules with complexities of $O(N^{2}D)$ or $O(ND^{2})$. In practical tests, the time taken for token selection is minimal, typically less than $1\%$ of the main computational cost. At cache steps, taking a caching ratio of $\mathcal{R}=90\%$ as an example, the computational cost of terms with a complexity of $O(ND^{2})$ is reduced to $10\%$ of the original, while the computational cost of terms with a complexity of $O(N^{2}D)$ is reduced to $1\%$. (However, as mentioned earlier, in practice, it is more efficient to shift all computations at cache steps to the MLP. Therefore, all terms with a complexity of $O(N^{2}D)$ at cache steps are ignored.)
Figure 12: A more detailed workflow for the proposed ToCa.The cache-and-reuse procedure is conducted on the model at all layers and timesteps.
A.5 Implemented Results on Class-Conditional Image Generation
Table 6: Quantitative comparison on class-to-image generation on ImageNet with 50 steps DDIM sampler as baseline on DiT-XL/2.
DiT-XL/2-G (cfg $=1.50$) 0.455 23.74 1.00$\times$ 4.40 2.43 0.80 0.59
$50\%$ steps 0.238 11.86 2.00$\times$ 4.74 3.18 0.79 0.58
$40\%$ steps 0.197 9.50 2.50$\times$ 5.15 3.81 0.78 0.57
$34\%$ steps 0.173 8.08 2.94$\times$ 5.76 4.58 0.77 0.56
$\textbf{FORA}(\mathcal{N}=2.5)$ 0.219 10.48 2.27$\times$ 6.59 3.83 0.79 0.55
$\textbf{FORA}(\mathcal{N}=3)$ 0.211 8.58 2.77$\times$ 6.43 3.88 0.79 0.54
ToCa ($\mathcal{N}=3,R=93\%$) 0.227 10.23 2.32$\times$ 4.74 3.04 0.80 0.57
ToCa ($\mathcal{N}=4,R=93\%$) 0.209 8.73 2.72$\times$ 5.11 3.60 0.79 0.56
In addition to the series of experiments conducted using the DDPM(Ho et al., 2020) sampling method on DiT, which have already been included in the main part, we also performed validation with the more practically relevant DDIM(Song et al., 2021) sampling method to further demonstrate the effectiveness of ToCa as shown in Table 6.
For instance, ToCa leads to 1.32 and 0.28 lower values in sFID and FID compared with FORA with a similar acceleration ratio of approximately 2.7×2.7$\times$2.7 ×, respectively, and achieves 0.21 lower values in FID compared to the method of directly reducing the sampling steps with an acceleration ratio of approximately 2.5×2.5$\times$2.5 ×. As a trade-off between acceleration and performance, we selected the scheme $\mathcal{N}=3,R=93\%$ as the final recommended approach for DDIM sampler.
A.6 Results on Higher-Resolution and more advanced text-to-image models
Table 7: Quantitative comparison in text-to-image generation for FLUX on Image Reward.
Method Latency(s) $\downarrow$ FLOPs(T) $\downarrow$ Speed $\uparrow$ Image Reward $\uparrow$
FLUX.1-dev (Labs, 2024) 33.85 3719.50 1.00$\times$ 1.202
$68\%$ steps 23.02 2529.26 1.47$\times$ 1.200
FORA (Selvaraju et al., 2024) 20.82 2483.32 1.51$\times$ 1.196
ToCa$(\mathcal{N}=2,R=90\%)$ 19.88 2458.06 1.51$\times$ 1.202
FLUX.1-schnell (Labs, 2024) 2.882 277.88 1.00$\times$ 1.133
$75\%$ steps 2.162 208.41 1.33$\times$ 1.139
$\textbf{FORA}^{1}$ (Selvaraju et al., 2024) 2.365 225.60 1.23$\times$ 1.129
$\textbf{FORA}^{2}$ (Selvaraju et al., 2024) 2.365 225.60 1.23$\times$ 1.124
$\textbf{FORA}^{3}$ (Selvaraju et al., 2024) 2.365 225.60 1.23$\times$ 1.123
ToCa$(\mathcal{N}=2,R=90\%)$ 1.890 181.30 1.53$\times$ 1.134
As shown in Table 7, we compared the performance of FORA and ToCa in generating high-resolution images ($1024\times 1024$) using the more advanced text-conditional image generation models FLUX.1-dev and FLUX.1-schnell(Labs, 2024). The former uses 50 sampling steps, while the latter, as a more efficient model, uses only 4 sampling steps. The evaluation of generation quality was conducted using Image Reward, a metric better suited to measuring human preference. The generated images were based on 1,632 prompts from the PartiPrompts(Yu et al., 2022) dataset to comprehensively evaluate the generation quality of the acceleration methods on both FLUX.1-dev and FLUX.1-schnell.
In this comparison, FORA (Selvaraju et al., 2024) represents acceleration on the FLUX.1-dev model with 50 sampling steps, where caching is performed every other step (i.e., $\mathcal{N}=2$). FORA1, FORA2, and FORA3 correspond to skipping the 2nd, 3rd, and 4th steps, respectively, during the 4-step generation process. ToCa demonstrated nearly lossless performance under a 1.5×1.5$\times$1.5 × acceleration, with Image Reward scores almost identical to the non-accelerated scenario in both the 50-step FLUX.1-dev and the 4-step FLUX.1-schnell models. In contrast, all configurations of FORA showed quality degradation. For example, in the 50-step FLUX.1-dev model, the 1.5×1.5$\times$1.5 × acceleration setting and in the 4-step FLUX.1-schnell model, a 1.2×1.2$\times$1.2 × acceleration setting both resulted in noticeable quality degradation. The corresponding visual results for FLUX.1-schnell are presented in Figure 13, demonstrating the lossless acceleration capability of ToCa.
Figure 13: Visualization examples for original FLUX.schnell(Labs, 2024) and proposed ToCa with almost lossless acceleration.
Figure 14: Computation ratio for different types of computation layers on different timesteps and layers on DiT. The dark blue lines correspond to fully computed fresh steps. (a) The computation ratio distribution of the Self-Attention layers. As mentioned earlier, performing partial computation on attention layers is less cost-effective compared to MLP layers. Therefore, we do not implement partial computation on attention layers; apart from fresh steps, all other steps directly reuse the corresponding cached features. (b) The computation ratio distribution of the MLP layers. As shown, the computation ratio increases with deeper layers and as the number of inferred timesteps during the inference phase grows.
A.7 Details for distribution figures
In this section, we provide detailed explanations of the various distribution plots mentioned in the main text as supplementary information.
Figure 2 illustrates the distribution of the Frobenius norm of the differences between the feature maps $x^{L}_{t}$ and $x^{L}_{t+1}$, which are the outputs of the last DiT Block at each timestep $t$ and the previous timestep $t+1$, respectively. The figure also presents the corresponding statistical frequency density of these Frobenius norm values for each token, based on 500 randomly generated samples produced by DiT. This analysis reveals the conclusion that different tokens exhibit varying levels of temporal redundancy.
Figure 2 illustrates the varying rates of error accumulation and propagation across different tokens. Specifically, Gaussian noise with an intensity of 0.5 was independently added to the $i_{th}$ token of the first layer at each step. The Frobenius norm was then computed between the output features of all tokens at the last layer of the same step and the corresponding features from the noise-free output. This process was repeated for all steps and all layers. Given that each noise propagation required re-running the inference process, a random subset of 100 samples from the DiT model was selected for this case study, and the noise propagation results were recorded for each iteration. This analysis led to the conclusion that different tokens exhibit varying rates of error accumulation and propagation.
Figure 5(a) illustrates the varying temporal redundancy across layers of different depths. For each timestep, the Frobenius norm of the differences between the features of the current timestep and the corresponding features of the previous timestep at a specific layer depth was computed for each token. The resulting Frobenius norm values were then used to plot the distribution alongside their corresponding statistical frequency densities. To clearly demonstrate the trends, we selected layers 1, 15, and 28 for visualization. The samples used were randomly chosen from 200 DiT samples.
Figure 5(b) shows the variation in the offset distribution of the output values from the last layer of a timestep when Gaussian noise is added to a single token at different layer depths within the same timestep. This is measured by adding Gaussian noise with an intensity of 0.5 to a single token in a specific layer at one timestep and comparing the deviation in the output features of the last layer at that timestep with the noise-free scenario. For clarity, this operation was performed on layers 1, 15, and 25 across all timesteps, using 200 randomly selected samples from the DiT model to generate the examples. It is worth noting that Figure 5(b) may appear at first glance to violate the normalization condition for frequency density distributions. This is due to the large variations in Frobenius norm values, which necessitated the use of a logarithmic scale on the horizontal axis. Figure 5(a) and (b) demonstrate two key conclusions: deeper layers exhibit poorer temporal redundancy, but errors introduced in deeper layers have a smaller impact on the output at the same timestep.
Figure 5(c) illustrates the results on the PixArt-$\alpha$ model by adding Gaussian noise with an intensity of $0.5\times\|x_{k}\|_{F}$ to a single token $x_{k}$ in the 10th layer (approximately the middle layer) at each timestep. This process was performed for three different types of layers (self-attention, cross-attention, and MLP). The Frobenius norm of the error induced by the noise was measured on the output of the last layer and normalized by the average Frobenius norm $\|x_{k}\|_{F}$ of tokens of the same type. The resulting distribution was plotted using 200 prompts randomly selected from the MS-COCO2017 dataset. It is important to note that the additional normalization step, based on the norm values of the tokens, was necessary because the norm values of tokens in self-attention, cross-attention, and MLP layers typically vary significantly. Normalization ensures a fair comparison across these layer types. Additionally, the distribution for the MLP layer in Figure 5(c) appears more dispersed. This is due to the generally larger variations in MLP output values across different prompts and timesteps. In practice, increasing the number of samples can make the distribution visually denser. However, given that each token requires a separate inference for the error propagation experiments, using 200 prompts already incurs a significant computational cost while remaining sufficient to reveal the trends.
Input: current timestep $t$, current layer id $l$.
Algorithm 1 ToCa
1:  if current timestep $t$ is a fresh step then
2:     Fully compute $\mathcal{F}^{l}(x)$.
3:     $\mathcal{C}^{l}(x):={\mathcal{F}^{l}}(x)$; # Update the cache.
4:  else
5:     $\mathcal{S}(x_{i})=\sum_{j=1}^{4}\lambda_{j}\cdot s_{j}$; # Compute the cache score for each token.
6:     $\mathcal{I}_{Compute}:=\text{TopK}(\mathcal{S}(x_{i}),R\%)$; # Fetch the index of computed tokens.
7:     for all tokens $x_{i}$ do
8:        if $i\in\mathcal{I}_{Compute}$ then
9:           Compute ${\mathcal{F}^{l}}(x_{i})$ through the neural layer.
10:           $\mathcal{C}^{l}(x_{i}):={\mathcal{F}^{l}}(x_{i})$; # Update the cache.
11:        end if
12:     end for
13:  end if
14:  return $\mathcal{F}^{l}(x)$. # return features for both cached and computed tokens for the next layer.
Figure 15: Pareto curve with FLOPs-FID to better evaluate the performance of ToCa on DiT with 50 ddim sampling steps as baseline.